Waterways Problem 05

The Vice President for Sales and Marketing at Waterways Corporation is planning for production needs to meet sales demand in the coming year. He is also trying to determine how the company’s profits might be increased in the coming year. This problem asks you to use cost-volume-profit concepts to help Waterways understand contribution margins of some of its products and decide whether to mass-produce any of them.

Waterways markets a simple water control and timer that it mass-produces. Last year, the company sold 751,000 units at an average selling price of $4.90 per unit. The variable costs were $2,575,930, and the fixed costs were $772,779.

What is the company’s break-even point in units and in dollars for this product?
Break-even point in units units
Break-even point in dollars $
What is the margin of safety, both in dollars and as a ratio? (Round ratio to 0 decimal places, e.g. 25%.)
Margin of safety in dollars $
Margin of safety ratio %
Answer #1

Selling price per unit = $4.90

Variable cost per unit = 2,575,930/751,000 = $3.43

Contribution margin per unit = 4.90 - 3.43 = 1.47

Fixed cost = $772,779

Breakeven sales in units = Fixed cost/ contribution margin

= 772,779/1.47

= 525,700 units

Breakeven sales in dollars = 525,700*4.90 = $2,575,930

Margin of safety sales = Total sales - Breakeven sales

= 751,000 - 525,700

= 225,300 units

Margin of safety sales in dollars

= 225,300*4.90

= $1,103,970

Margin of safety ratio

= (margin of safety sales/ total sales)*100

= (225,300/751,000)*100

= 30%

If you find the answer helpful please upvote.